find the altitude of an isosceles triangle with sides of 13 inches and a base of 10 inches
Answer by stanbon(75887)

find the altitude of an isosceles triangle with sides of 13 inches and a base of 10 inches.
----------
Draw the triangle as described.
Draw the altitude from the base to the vertex opposite.
The altitude bisects the base into two segments of size 5.
Use Pythagoras to find the altitude.
13^2= (altitude)^2 + 5^2
(alt)^2=169-25=144
alt=12
Cheers,
Stan H.